# The Latency Price of Threshold Cryptosystem in Blockchains

### v1

@@ 6.3 Proof of Correctness
Let $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ denote the protocol. The proofs of Agreement, Termination, Validity and Total Order properties are identical to that of Theorem 5.1. Thus, we focus on the Secrecy property.
@@
We first prove that, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$. For the sake of contradiction, suppose that for some $r\in\mathbb{N}$, an honest party outputs $(r,m,\sigma^{\prime})$ where $\sigma^{\prime}\neq{\sf Eval}(s,(r,m))$. By the protocol, the honest party outputs ${\sf finalize}(r,m)$ in $\sf MBB_{FT}$. Also, as per the protocol and the Robustness property of $\sf TC$, $\sigma^{\prime}={\sf Eval}(s,(r,m^{\prime}))$ for some $m^{\prime}\neq m$. There are two cases:
@@
Second, if $\sigma^{\prime}$ is obtained from the fast path, then by the Unpredictability property of $\sf TC$, at least $t_{\sf sec}^{\prime}-|\mathcal{C}|$ honest parties reveal their ${\sf PEval}(\llbracket{s}\rrbracket^{\prime},(r,m^{\prime}))$ for the fast path. According to the protocol, these honest parties call ${\sf prefinalize}(r,m^{\prime})$ in $\sf MBB_{FT}$. Since $t_{\sf fin}=t_{\sf sec}^{\prime}$, the message $m^{\prime}$ is globally finalized since $\sf MBB_{FT}$ has the finalization threshold $t_{\sf fin}$. Again, the Agreement property of $\sf MBB_{FT}$ is violated, contradiction. Therefore, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$.
@@
Theorem 6.3
Algorithm 3 achieves ${\sf GFT}_{r}={\sf GRT}_{r}$ for round $r$ if the execution is synchronous for $r$.
@@
Consider any synchronous execution. By Definition 15, all honest parties prefinalize the same message $m$ for round $r$ at the same physical time ${\sf GFT}_{r}$. Thus, as per  Algorithm 3, all honest parties reveal their $\sf TC$ output shares for $(r,m)$ at time ${\sf GFT}_{r}$ in the fast path, which allows the adversary to reconstruct the $\sf TC$ output at time ${\sf GFT}_{r}$, which implies ${\sf GRT}_{r}\leq{\sf GFT}_{r}$. By Lemma 1, ${\sf GRT}_{r}\geq{\sf GFT}_{r}$. Therefore, we conclude that ${\sf GFT}_{r}={\sf GRT}_{r}$ for any $r$ in any synchronous execution. ∎
Theorem 6.4
@@
Consider any optimistic execution and any round $r\in\mathbb{N}$. As per Algorithm 3, honest parties prefinalize a message and send its $\sf TC$ output share for the fast path simultaneously. Recall from Definition 16, in optimistic executions, all messages have the same delay, and all parties are honest and finalize the same message at the same physical time. Therefore, all honest parties locally receive $t_{\sf fin}$ ${\tt PREFIN}$ messages and $t_{\sf rec}$ $\sf TC$ output shares for $(r,m)$ simultaneously. This implies that the local finalization and reconstruction occur simultaneously. Then, by a similar induction argument as in Theorem 5.3, ${\sf L}_{r}=0$ for any round $r\in\mathbb{N}$ in any optimistic execution. ∎
In practice, different honest parties may prefinalize at different physical times due to a lack of synchrony. Moreover, honest parties may need to wait slightly longer after local finalization to receive additional shares from the fast path since the reconstruction threshold of the fast path is higher than the finalization threshold. Despite this, the latency of Algorithm 3 remains significantly lower than one message delay. The evaluation in Section 7.3 demonstrates that the fast path reduces the latency overhead by $71\%$ compared to the slow path (Algorithm 1), which has a latency of one message delay.
7 Distributed Randomness: A Case Study
In this section, we implement and evaluate distributed randomness as a concrete example of blockchain-native threshold cryptosystem, to demonstrate the effectiveness of our solution in reducing the latency for real-world blockchains. We implement the fast path (Algorithm 3) for Das et al. [23], which is a distributed randomness scheme designed for proof-of-stake blockchains and is deployed in the Aptos blockchain [10]. We then compare our latency (using both micro-benchmarks and end-to-end evaluation) with the [23], that implements the strawman protocol (Algorithm 1).
In the rest of the section, we first provide the brief overview of [23] necessary to describe the implementation of our fast-path protocol atop their scheme.
7.1 Overview of Das et al. [23]
@@ 7.2 Implementation
Consensus. Here we explain how the consensus in [23] satisfies the definition of $\sf MBB_{FT}$ (Definition 5). Das et al. [23] is built on the Aptos blockchain, which uses an improved version  [^7] At of June 2024, the consensus latency of Jolteon on the Aptos mainnet is $4$ message delays, improved from $5$ message delays [29]. Further improvements to $3$ message delay are currently in progress [9]. of Jolteon [29] as its consensus protocol. Jolteon tolerates an adversary capable of corrupting parties holding up to $t$ stakes out of a total of $n=3t+1$ stakes under partial synchrony. In Jolteon, each message $m$ in $\sf MBB_{FT}$ is a block containing transactions, and only one party (called leader) can propose a block for each round.
  • Prefinalization. A party at local round $\leq r$ calls ${\sf prefinalize}(r,m)$ upon receiving a quorum certificate (QC) for $m$ of round $r$, where a QC consists of votes from parties with combined stakes of $2/3$-rd of the total stakes. When a party prefinalizes a block $m$ of round $r$, it also implicitly prefinalizes all previous rounds that are not yet finalized (either parent blocks or $\bot$). In Jolteon, each party votes for at most one block per round, so that at most one QC can be formed for each round by quorum intersection. This means that for each round, all honest parties can only prefinalize at most one block.
  • Finalization. A block $m$ is globally finalized for round $r$ if and only if parties with combined $t_{\sf fin}=2t+1$ stakes (or $t_{\sf fin}-t=t+1$ honest stakes) calls ${\sf prefinalize}(r,m)$. A party locally finalizes $m$ for round $r$ once it receives $({\tt PREFIN},r,m)$ messages from parties with $t_{\sf fin}$ stakes, which also finalizes all previous rounds with parent blocks or $\bot$.
Rounding. Recall that the fast path (Algorithm 3) requires sharing the same secret two sets of thresholds, i.e., $t_{\sf sec}<t_{\sf rec}$ for the slow path and $t_{\sf sec}^{\prime}<t_{\sf rec}^{\prime}$ for the fast path. Consequently, we augment the rounding algorithm of [23] to additionally take $(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$ as input, and output the weight threshold $w^{\prime}$ for the fast path. That is, ${\sf Round^{\prime}}(\bm{S},t_{\sf sec},t_{\sf rec},t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})\rightarrow(\bm{W},w,w^{\prime})$ where $w,w^{\prime}$ are the weight thresholds of the slow path and fast path, respectively. ${\sf Round^{\prime}}$ also provides the same guarantees for the thresholds of fast path, namely any subset of parties with a combined stake $<t_{\sf sec}^{\prime}$ will always have a combined weight $<w^{\prime}$, and any subset of parties with a combined stake $\geq t_{\sf rec}^{\prime}$ will always have a combined weight $\geq w^{\prime}$. As in Algorithm 3, $t_{\sf sec}^{\prime}$ is set to be equal to the consensus finalization threshold $t_{\sf fin}$, and $t_{\sf rec}^{\prime}$ is set to be $n$ (total stake).
@@
Randomness generation. As described in Algorithm 3, parties reveal their VUF shares ($\sf TC$ output shares) for the fast path upon prefinalizing a block, and for the slow path upon finalizing a block. For both paths, the parties collect the VUF shares and are ready to execute the block as soon as the randomness ($\sf TC$ output) is reconstructed from either path. As in [23], parties use the VUF output evaluated on the secret $s$ and the round (and epoch) number of the finalized block as the $\sf TC$ output  [^8] In the implementation of Algorithm 3 with Jolteon as the $\sf MBB$ protocol, it is safe to omit $m$ in the evaluation function, since all honest parties always prefinalize the same block for each round as mentioned in section consensus.. Since the parties are weighted, similar to Section 7.1, we implement the weighted extension of Algorithm 3.
Implementation. We implement fast-path in Rust, atop the open-source Das et al. [23] implementation [10] on the Aptos blockchain. Our implementation [11] is also deployed in Aptos blockchain. For communication, we use the ${\tt Tokio}$ library [2]. For cryptography, we use the ${\tt blstrs}$ library [1], which implements efficient finite field and elliptic curve arithmetic. Similar to Das et al. [23], our implementation runs the share verification step of different parties in parallel and parallelizes the VUF derivation using multi-threading.
7.3 Evaluation Setup

### v3

@@ 0.C.1 Analysis of Algorithm 2
Let $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ denote the protocol.
@@
We first prove that, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$. By the Agreement and Termination properties of $\sf MBB_{FT}$, all honest parties output ${\sf finalize}(r,m)$ in $\sf MBB_{FT}$. By Definition 4,
  1. 1.
    There exist $t_{\sf fin}$ parties (or equivalently $t_{\sf fin}-|\mathcal{C}|$ honest parties) that have sent $({\tt PREFIN},r,m)$.
  2. 2.
    If $m\neq\bot$, then no honest party sent $({\tt PREFIN},r,m^{\prime})$ such that $m^{\prime}\in{\mathcal{M}}$ and $m^{\prime}\neq m$.
  3. 3.
    If $m=\bot$, then there exists at most one $m^{\prime}\in{\mathcal{M}}$ such that any honest party may send $({\tt PREFIN},r,m^{\prime})$.
For the sake of contradiction, suppose that for some $r\in\mathbb{N}$, an honest party $h$ outputs $(r,m,\sigma^{\prime})$ where $\sigma^{\prime}\neq{\sf Eval}(s,(r,m))$. As per the protocol and the Robustness property of $\sf TC$, $\sigma^{\prime}={\sf Eval}(s,(r,m^{\prime}))$ for some $m^{\prime}\neq m$. Note that in Algorithm 2 an honest party reveals its $\sf TC$ output share for any $(r,m^{\prime})$ only upon prefinalizing $(r,m^{\prime})$ or finalizing $(r,m^{\prime})$. This implies, by the Unpredictability property of $\sf TC$, that at least $t_{\sf sec}-|\mathcal{C}|$ honest parties have revealed their $\sf TC$ output shares ${\sf PEval}(\llbracket{s}\rrbracket,(r,m^{\prime}))$. Let $T_{m^{\prime}}$ be the set indicating these honest parties.
  • If any honest party $i\in T_{m^{\prime}}$ finalizes a message $(r,m^{\prime})$ for $m^{\prime}\neq m$, then this violates the Agreement property of $\sf MBB_{FT}$, and hence a contradiction.
  • This implies that all parties in $T_{m^{\prime}}$ has sent $({\tt PREFIN},r,m^{\prime})$ in $\sf MBB_{FT}$. If $m\neq\bot$, then no honest party should have sent $({\tt PREFIN},r,m^{\prime})$ such that $m^{\prime}\in{\mathcal{M}}$ and $m^{\prime}\neq m$, contradiction. If $m=\bot$, then all $|T_{m^{\prime}}|\geq t_{\sf sec}-|\mathcal{C}|=t_{\sf fin}-|\mathcal{C}|$ honest parties in $T_{m^{\prime}}$ has sent $({\tt PREFIN},r,m^{\prime})$ in $\sf MBB_{FT}$. This implies that the message $m^{\prime}$ is globally finalized, again violating the Agreement property of $\sf MBB_{FT}$, hence a contradiction.
Therefore, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$.
For any corrupted party, the same argument above applies; thus the adversary cannot learn ${\sf Eval}(s,(r,m^{\prime}))$ where $m^{\prime}\neq m$.
Agreement.
Termination.
Validity.
Total Order.
Theorem 0.C.2
Proof
Recall from Section 2, for any round $r$, ${\sf GFT}_{r}$ and ${\sf GRT}_{r}$ are the global finalization time and global reconstruction time of round $r$, respectively. Consider any execution. By the definition of ${\sf GFT}_{r}$, $t_{\sf fin}-|\mathcal{C}|$ honest parties have sent $({\tt PREFIN},r,m)$ at time ${\sf GFT}_{r}$. According to Algorithm 2, these honest parties have also revealed their $\sf TC$ output shares for $(r,m)$ at time ${\sf GFT}_{r}$. Since, $t_{\sf rec}=t_{\sf fin}$, this implies that the adversary can adversary to reconstruct the $\sf TC$ output at time ${\sf GFT}_{r}$, and hence ${\sf GRT}_{r}\leq{\sf GFT}_{r}$. From Lemma 1, we have that ${\sf GRT}_{r}\geq{\sf GFT}_{r}$. Therefore, we conclude that ${\sf GFT}_{r}={\sf GRT}_{r}$ for any $r$ in any execution.
Theorem 0.C.3
Proof
Appendix 0.D Ramp Blockchain-Native Threshold Cryptography
0.D.1 Impossibility
Proof (Proof of Theorem 5.1)
For the sake of contradiction, suppose that there exists a protocol $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ with ramp thresholds, such that for any given round $r$, in all executions we have ${\sf GRT}_{r}={\sf GFT}_{r}$. Let $\mathcal{E}$ be one such execution and let $\tau_{\mathcal{E}}={\sf GRT}_{r}={\sf GFT}_{r}$. Let $s$ denote the secret shared between the parties. Since $\sf MBB_{FT}$ has the finalization threshold $t_{\sf fin}$, at least $t_{\sf fin}-|\mathcal{C}|$ honest parties have sent $({\tt PREFIN},r,m)$ at $\tau_{\mathcal{E}}$. Similarly, by the definition of ${\sf GRT}_{r}$, at least $t_{\sf rec}-|\mathcal{C}|$ honest parties have revealed their $\sf TC$ output shares ${\sf PEval}(\llbracket{s}\rrbracket,(r,m))$ at $\tau_{\mathcal{E}}$. Without loss of generality, we can assume that exactly $t_{\sf fin}-|\mathcal{C}|$ honest parties have sent $({\tt PREFIN},r,m)$ and $t_{\sf rec}-|\mathcal{C}|$ honest parties have revealed their $\sf TC$ output shares at $\tau_{\mathcal{E}}$. Let $h$ be any honest party that sent $({\tt PREFIN},r,m)$ at time $\tau_{\mathcal{E}}$ in the execution $\mathcal{E}$. There are two cases, which we denote with (1) and (2) below.
  1. (1)
    $h$ also reveals its $\sf TC$ output share ${\sf PEval}(\llbracket{s}\rrbracket_{h},(r,m))$ at time $\tau_{\mathcal{E}}$. Consider another execution $\mathcal{E}^{\prime}$ identical to $\mathcal{E}$ up to time $\tau_{\mathcal{E}}$ with the only difference that, $h$ sends $({\tt PREFIN},r,m)$ and reveals ${\sf PEval}(\llbracket{s}\rrbracket_{h},(r,m))$ at time time $\tau_{\mathcal{E}}+\epsilon$ for some $\epsilon>0$ (say due to asynchrony in computation or communication). Therefore, $\tau_{\mathcal{E}}+\epsilon$ is the global finalization time of execution $\mathcal{E}^{\prime}$.
    Now, consider the time $\tau_{\mathcal{E}}$ in the new execution $\mathcal{E}^{\prime}$. The message $m$ is not globally finalized at $\tau_{\mathcal{E}}$ for $r$ since only $t_{\sf fin}-|\mathcal{C}|-1$ honest parties have sent $({\tt PREFIN},r,m)$. However, since $t_{\sf rec}>t_{\sf sec}$, there are $t_{\sf rec}-|\mathcal{C}|-1\geq t_{\sf sec}-|\mathcal{C}|$ honest parties that have revealed their $\sf TC$ output shares ${\sf PEval}(\llbracket{s}\rrbracket,(r,m))$. This implies that in execution $\mathcal{E}^{\prime}$ the adversary $\mathcal{A}$ can compute the $\sf TC$ output ${\sf Eval}(s,(m,r))$ at $\tau_{\mathcal{E}}$, which is earlier than its global finalization time $\tau_{\mathcal{E}}+\epsilon$. This contradicts Lemma 1.
  2. (2)
    $h$ does not reveal its $\sf TC$ output share at time $\tau_{\mathcal{E}}$. Similarly, consider another execution $\mathcal{E}^{\prime}$ identical to $\mathcal{E}$ up to time $\tau_{\mathcal{E}}$ with the only difference that, $h$ sends $({\tt PREFIN},r,m)$ at time $\tau_{\mathcal{E}}+\epsilon$ for some $\epsilon>0$ (say due to asynchrony in computation or communication). Therefore, for the new execution $\mathcal{E}^{\prime}$, $\tau_{\mathcal{E}}+\epsilon$ is the global finalization time, and the global reconstruction time remains $\tau_{\mathcal{E}}$. Then, for the new execution $\mathcal{E}^{\prime}$, the global reconstruction comes before the global finalization, contradicting Lemma 1. ∎
0.D.2 Analysis of Algorithm 3
Theorem 0.D.1
Proof
Let $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ denote the protocol. The proofs of Agreement, Termination, Validity and Total Order properties are identical to that of Theorem 0.C.1. Thus, we focus on the Secrecy property.
Secrecy.
We first prove that, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$. By the Agreement and Termination properties of $\sf MBB_{FT}$, all honest parties output ${\sf finalize}(r,m)$ in $\sf MBB_{FT}$. By Definition 4,
  1. 1.
    There exist $t_{\sf fin}$ parties (or equivalently $t_{\sf fin}-|\mathcal{C}|$ honest parties) that have sent $({\tt PREFIN},r,m)$.
  2. 2.
    If $m\neq\bot$, then no honest party sent $({\tt PREFIN},r,m^{\prime})$ such that $m^{\prime}\in{\mathcal{M}}$ and $m^{\prime}\neq m$.
  3. 3.
    If $m=\bot$, then there exists at most one $m^{\prime}\in{\mathcal{M}}$ such that any honest party may send $({\tt PREFIN},r,m^{\prime})$.
For the sake of contradiction, suppose that for some $r\in\mathbb{N}$, an honest party $h$ outputs $(r,m,\sigma^{\prime})$ where $\sigma^{\prime}\neq{\sf Eval}(s,(r,m))$. As per the protocol and the Robustness property of $\sf TC$, $\sigma^{\prime}={\sf Eval}(s,(r,m^{\prime}))$ for some $m^{\prime}\neq m$. Also, as per the protocol and the Robustness property of $\sf TC$, $\sigma^{\prime}={\sf Eval}(s,(r,m^{\prime}))$ for some $m^{\prime}\neq m$. There are two cases:
@@ 0.D.2 Analysis of Algorithm 3
Second, if $\sigma^{\prime}$ is obtained from the fast path, then by the Unpredictability property of $\sf TC$, at least $t_{\sf sec}^{\prime}-|\mathcal{C}|$ honest parties reveal their ${\sf PEval}(\llbracket{s}\rrbracket^{\prime},(r,m^{\prime}))$ for the fast path. If $m\neq\bot$, then no honest party should have sent $({\tt PREFIN},r,m^{\prime})$ such that $m^{\prime}\in{\mathcal{M}}$ and $m^{\prime}\neq m$, contradiction. If $m=\bot$, then at least $t_{\sf sec}^{\prime}-|\mathcal{C}|=t_{\sf fin}-|\mathcal{C}|$ honest parties has sent $({\tt PREFIN},r,m^{\prime})$ in $\sf MBB_{FT}$. This implies that the message $m^{\prime}$ is globally finalized, again violating the Agreement property of $\sf MBB_{FT}$, hence a contradiction. Therefore, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$.
@@
Theorem 0.D.2
Algorithm 3 achieves ${\sf GFT}_{r}={\sf GRT}_{r}$ for round $r$ if the execution is synchronized for $r$.
@@
Consider any synchronized execution. By the definition of synchronized execution, all honest parties prefinalize the same message $m$ for round $r$ at the same physical time ${\sf GFT}_{r}$. Thus, as per  Algorithm 3, all honest parties reveal their $\sf TC$ output shares for $(r,m)$ at time ${\sf GFT}_{r}$ in the fast path, which allows the adversary to reconstruct the $\sf TC$ output at time ${\sf GFT}_{r}$, which implies ${\sf GRT}_{r}\leq{\sf GFT}_{r}$. By Lemma 1, ${\sf GRT}_{r}\geq{\sf GFT}_{r}$. Therefore, we conclude that ${\sf GFT}_{r}={\sf GRT}_{r}$ for any $r$ in any synchronized execution. ∎
Theorem 0.D.3
@@
Consider any optimistic execution and any round $r\in\mathbb{N}$. As per Algorithm 3, honest parties prefinalize a message and send its $\sf TC$ output share for the fast path simultaneously. Recall from the definition of optimistic execution, in optimistic executions, all messages have the same delay, and all parties are honest and finalize the same message at the same physical time. Therefore, all honest parties locally receive $t_{\sf fin}$ ${\tt PREFIN}$ messages and $t_{\sf rec}$ $\sf TC$ output shares for $(r,m)$ simultaneously. This implies that the local finalization and reconstruction occur simultaneously. Then, by a similar induction argument as in Theorem 0.C.3, ${\sf L}_{r}=0$ for any round $r\in\mathbb{N}$ in any optimistic execution. ∎
In practice, different honest parties may prefinalize at different physical times due to a lack of synchrony. Moreover, honest parties may need to wait slightly longer after local finalization to receive additional shares from the fast path since the reconstruction threshold of the fast path is higher than the finalization threshold. Despite this, the latency of Algorithm 3 remains significantly lower than one message delay. The evaluation in Section 0.E.3 demonstrates that the fast path reduces the latency overhead by $71\%$ compared to the slow path (Algorithm 1), which has a latency of one message delay.
Appendix 0.E Distributed Randomness: A Case Study
0.E.1 Overview of Das et al. [23]
@@ 0.E.2 Implementation
Consensus. Here we explain how the consensus in [23] satisfies the definition of $\sf MBB_{FT}$ (Definition 4). Das et al. [23] is built on the Aptos blockchain, which uses a latency-improved version [10] of Jolteon [29] as its consensus protocol. Jolteon tolerates an adversary capable of corrupting parties holding up to $t$ stakes out of a total of $n=3t+1$ stakes under partial synchrony. In Jolteon, each message $m$ in $\sf MBB_{FT}$ is a block containing transactions, and only one party (called leader) can propose a block for each round.
  • Prefinalization. A party at local round $\leq r$ sends $({\tt PREFIN},r,m)$ upon receiving a quorum certificate (QC) for $m$ of round $r$, where a QC consists of votes from parties with combined stakes of $2/3$ of the total stakes. When a party prefinalizes a block $m$ of round $r$, it also implicitly prefinalizes all previous rounds that are not yet finalized (either parent blocks or $\bot$). In Jolteon, each party votes for at most one block per round, so that at most one QC can be formed for each round by quorum intersection. This means that for each round, all honest parties can only prefinalize at most one block.
  • Finalization. A block $m$ is globally finalized for round $r$ if and only if parties with combined $t_{\sf fin}=2t+1$ stakes (or $t_{\sf fin}-t=t+1$ honest stakes) send $({\tt PREFIN},r,m)$. A party locally finalizes $m$ for round $r$ once it receives $({\tt PREFIN},r,m)$ messages from parties with $t_{\sf fin}$ stakes, which also finalizes all previous rounds with parent blocks or $\bot$.
Rounding. Recall that the fast path (Algorithm 3) requires sharing the same secret two sets of thresholds, i.e., $t_{\sf sec}<t_{\sf rec}$ for the slow path and $t_{\sf sec}^{\prime}<t_{\sf rec}^{\prime}$ for the fast path. Consequently, we augment the rounding algorithm of [23] to additionally take $t_{\sf sec}^{\prime},t_{\sf rec}^{\prime}$ as input, and output the weight threshold $w^{\prime}$ for the fast path. That is, ${\sf Round^{\prime}}(\bm{S},t_{\sf sec},t_{\sf rec},t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})\rightarrow(\bm{W},w,w^{\prime})$ where $w,w^{\prime}$ are the weight thresholds of the slow path and fast path, respectively. ${\sf Round^{\prime}}$ also provides the same guarantees for the thresholds of fast path, namely any subset of parties with a combined stake $<t_{\sf sec}^{\prime}$ will always have a combined weight $<w^{\prime}$, and any subset of parties with a combined stake $\geq t_{\sf rec}^{\prime}$ will always have a combined weight $\geq w^{\prime}$. As in Algorithm 3, $t_{\sf sec}^{\prime}=t_{\sf fin},t_{\sf rec}^{\prime}=\min(t_{\sf sec}^{\prime}+(t_{\sf rec}-t_{\sf sec}),n)$.
@@
Randomness generation. As described in Algorithm 3, parties reveal their VUF shares ($\sf TC$ output shares) for the fast path upon prefinalizing a block, and for the slow path upon finalizing a block. For both paths, the parties collect the VUF shares and are ready to execute the block as soon as the randomness ($\sf TC$ output) is reconstructed from either path. As in [23], parties use the VUF output evaluated on the secret $s$ and the round (and epoch) number of the finalized block as the $\sf TC$ output  [^7] In the implementation of Algorithm 3 with Jolteon as the $\sf MBB$ protocol, it is safe to omit $m$ in the evaluation function, since all honest parties always prefinalize the same block for each round as mentioned in section consensus.. Since the parties are weighted, similar to Section 0.E.1, we implement the weighted extension of Algorithm 3.
Implementation Details. We implement fast-path in Rust, atop the open-source Das et al. [23] implementation [11] on the Aptos blockchain. We worked with Aptos to deploy our protocol on the Aptos blockchain. For communication, we use the ${\tt Tokio}$ library [2]. For cryptography, we use the ${\tt blstrs}$ library [1], which implements efficient finite field and elliptic curve arithmetic. Similar to Das et al. [23], our implementation runs the share verification step of different parties in parallel and parallelizes the VUF derivation using multi-threading.
0.E.3 Evaluation Setup